Manuel Kauers Hagenberg, 27. Juni 2003 RISC-Linz

Werden

• Here are, first of all, the forms:

ich  du  er
Präsens  werde  wirst  wird
Präteritum wurde  wurdest  wurde
Perfekt  bin geworden bist geworden ist geworden
wir  ihr  sie
Präsens  werden  werdet  werden
Präteritum wurden  wurdet  wurden
Perfekt  sind geworden seit geworden sind geworden

• The simple meaning of “werden” is “to become”.

• Examples

Ich werde verrückt!  I become crazy!
Du bist aber groß geworden! You have grown a lot!
Er ist Professor geworden  He has become a professor
Er wurde Professor  He became a professor

• Besides “to become”, “werden” is used for two other purposes: the future and the passive.

• The future is built by a conjugated form of “werden” plus an infinitive:

future = werden + infinitive

• Examples

Ich werde morgen nicht kommen  I won’t come tomorrow
Wir werden alle irgendwann sterben We will all die somewhen
Morgen wird es regnen  It will rain tomorrow

• The passive is built by a conjugated form of “werden” plus a PPP:

passive = werden + PPP

• Examples Du wirst überschätzt. You are overestimated. Das Auto wird verkauft. The car is sold. Wir werden von Frau Mayr zum Kaffee eingeladen.

We are invited to coffee by Frau Mayr.

• Passive in Pretäritum works by using the respective form of werden: Das Auto wurde verkauft, Wir wurden eingeladen,… • Passive in Perfect works by using a special PPP of werden: worden (i.e.

the ge- is omitted): Das Auto ist verkauft worden, Wir sind eingeladen worden,…

• Passive Future works by straight forward application of the rules: Das Auto wird verkauft werden, Wir werden eingeladen werden,…

• Frequently used to avoid inpersonal passive is the word man:

Es wurde viel getrunken – Much has been drunken vs. Man hat viel getrunken – One has drunken much

Exercise

• Was wirst Du im Sommer machen? (Write a few sentences about your plans, using the future forms)

• Convert the following sentences into passive. Example: Ich verkaufe das Auto → Das Auto wird von mir verkauft

Hans ißt jeden Tag eine Tafel Schokolade. 1492 entdeckte Kolumbus Amerika. 1492 hat Kolumbus Amerika entdeckt. Peter nimmt Susanne mit zur Uni.

• Verb forms can be viewed as elements of five dimensional space

Person × Numerus × Modus × Tempus × Genus

where

Person = {1., 2., 3.}
Numerus = {Singular, Plural}
Modus = {Indikativ, Konjunktiv}
Tempus = {Perfekt, Präteritum, Präsens, Futur, … }
Genus = {Aktiv, Passiv}

Take a regular word of your choice and write down all the forms of it. (Leave out the Konjunktiv form which we didn’t have so far.)